OpenStudy (anonymous):

What is the equation of the line passing through (–3, 6) and parallel to the line y = 4x – 2 in standard form?

OpenStudy (saifoo.khan):

pretriceing??/

OpenStudy (anonymous):

passing or pa$$ing?

OpenStudy (saifoo.khan):

prestressing preprocessing perthracing preracing ??

OpenStudy (anonymous):

I got the same thing when I tried to say c-o-m-p-a-s-s-i-o-n they wont let u type is the a$$ word

OpenStudy (anonymous):

she/he means p-a-s-s-i-n-g

OpenStudy (anonymous):

I think :)

OpenStudy (saifoo.khan):

LOL.

OpenStudy (anonymous):

anyways, now that that's established- since it's parallel, keep the slope, make the y intercept b, plug in your y and x values, and solve for b.

OpenStudy (anonymous):

y=4x+b 6=4(-3)+b 6=-12+b 18=b

OpenStudy (anonymous):

so ur line is y=4x+18

OpenStudy (anonymous):

Is Dat Equal To 4x + y = 18 ?